# High-precision direct decay energy measurements of the electron-capture decay of 97Tc

### v1

@@ Abstract
A direct measurement of the ground-state-to-ground-state electron-capture decay () value of 97Tc has been conducted employing the high resolving power phase-imaging ion-cyclotron-resonance technique with the double Penning trap mass spectrometer JYFLTRAP. The resulting value for 97Tc is 324.82(21) keV, exhibiting a precision approximately 19 times higher than the value adopted in the newest Atomic Mass Evaluation (AME2020) and differing by 1.2. Furthermore, by combining this refined value with nuclear energy-level data for the decay-daughter 97Mo, a potential ultra-low Q-value transition, possibly of allowed type, 97Tc (9/2+, ground state) 97Mo∗ (320(1) keV), was evaluated for future long-term neutrino-mass determination experiments. The ground-state-to-excited-state electron-capture decay value () of this transition was determined to be 4.8(10) keV, confirming it to be energetically allowed with a confidence level of exceeding 4. The captures of electrons occupying the L and higher shells for this transition are energetically allowed, giving a value of 2.0(10) keV for the closest distance of to the allowed binding energy of the L1 shell. To predict partial half-lives and energy-release distributions for this transition, the atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model have been employed. Dominant correction terms such as exchange and overlap corrections, as well as shake-up and shake-off effects, were included in the final results. Moreover, the normalized distribution of released energy in the electron-capture decay of 97Tc to excited states of 97Mo, is compared with that of 163Ho, which is being used for electron-neutrino-mass determination. A pseudo-experiment technique was introduced to calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions.
@@ I Introduction
Neutrinos are electrically neutral fermions that interact only via the weak force and gravity, and they have an incredibly tiny rest mass compared to other elementary particles. Neutrinos do not participate in electromagnetic or strong interactions, thus allowing them to pass through matter unimpeded. Up to now, three neutrino flavors have been found: electron neutrino (), muon neutrino (), and tau neutrino (). For each neutrino, there also exists a corresponding antineutrino, which also has spin of 1/2 and no electric charge Petcov (2013). Although neutrinos were initially thought to be massless as estimated by the standard model (SM), the three neutrino mass eigenstates do not show a one-to-one correspondence to the three flavors, each flavor being a quantum superposition of all three mass states Lesgourgues and Pastor (2006). Neutrinos oscillate between flavors, making their detection even more intriguing Fukuda et al. (1998); SNO Collaboration (2002); Gerbino and Lattanzi (2018). Neutrino-oscillation experiments reveal that at least two neutrino-mass eigenstates have a non-zero rest mass, leading to a mass difference between the three mass eigenstates. Laboratory experiments and cosmological observations have provided experimental data on differences in the squares of the neutrino masses, an upper limit on their sum, and an upper limit on the mass of the electron (anti)neutrino. Among these experiments, the most direct and model-independent way to pin down the exact mass value of the electron antineutrino (m) is to measure the electron energy spectrum of single decay, e.g., the Karlsruhe Tritium Neutrino (KATRIN) experiment Drexlin et al. (2013); Aker et al. (2019, 2022), which has recently pushed the absolute electron antineutrino mass scale boundary to 0.45 eV/c2 at the 90% confidence level (CL) Aker et al. (2024). KATRIN will finally set the m upper limit with a sensitivity of 0.2 eV/c2 at 90% CL. Other single--decay experiments, Project 8 and PTOLEMY project, using cyclotron radiation emission spectroscopy and the development of atomic tritium sources, will eventually allow a sensitivity of tritium-based neutrino-mass experiments to m beyond the design sensitivity of KATRIN Esfahani et al. (2017); Betti et al. (2019) down to 0.04 eV/c2. An alternative method, employing electron capture in 163Ho (ECHo), an experiment to calorimetrically measure the electron-capture (EC) deexcitation spectrum Gastaldo et al. (2014, 2017); Velte et al. (2019), has set a current limit of 150 eV/c2 for the electron neutrino mass Velte et al. (2019).
@@
In single--decay experiments aimed at determining the mass of the electron antineutrino, a value of the decay as small as possible is sought. A smaller value leads to a higher fraction of effective decays within a given energy interval near the endpoint region. Currently, the running direct (anti)neutrino-mass-determination experiments rely exclusively on ground-state-to-ground-state (gs-to-gs) decay cases, such as 3H ( decay) and 163Ho (EC). However, ongoing intensive searches are actively exploring isotopes having low -value ground-state-to-excited-state (gs-to-es) /EC decays Haaranen and Suhonen (2013); Suhonen (2014); Ejiri et al. (2019) using high-precision Penning-trap mass spectrometry (PTMS) including JYFLTRAP, LEBIT, CPT, ISOTRAP, and SHIPTRAP Penning traps Sandler et al. (2019); Karthein et al. (2019); De Roubin et al. (2020); Ge et al. (2021a, b, 2022a); Eronen et al. (2022); Ge et al. (2022a, b); Ramalho et al. (2022); Gamage et al. (2022); Keblbeck et al. (2023); Ge et al. (2023); Valverde et al. (2024); Ge et al. (2024a, b). PTMS is the most precise and accurate technique for determining atomic masses and values. It remains the only direct method capable of measuring decay values to sub-keV precision or better, allowing verification of potential ultra-low (< 1 keV) -value transitions.
@@ II Experimental method
To minimize any shifts in the ratio of the 97Tc+-97Mo+ pair, the positions of the phase spots for the magnetron and reduced cyclotron motions were carefully chosen to maintain the angle within a few degrees. This was done to account for the conversion of the reduced cyclotron motion to the magnetron motion and the potential distortion of the ion-motion projection onto the detector, ensuring that the shift remained well below 10-10 Eliseev et al. (2014). Figure 2 shows a representative measurement with "reduced cyclotron" and "magnetron" phase spots relative to the center spot. During the measurements, the excitation delay of was systematically scanned over one magnetron period, while the extraction delay varied over one cyclotron period, accounting for any residual motion that might have affected the spots. The interleaved measurements of for 97Tc+-97Mo+ ions had a total data accumulation time of approximately 4.9 hours.
@@
Final state | Decay type | (Lit.) | (T. W.) | (Lit.) | (T. W.) | (T. W.) | (T. W.) | (T. W.) | (T. W.) | |
97Mo () | {allowed} | 320.0(40) | 324.82(21) | 0.0(41) | 4.8(10) | 320.0(10) | 4.8 | 2.0(10) | 2.2(10) | 4.3(10) |
@@ III Results and discussion
From the measurements of the cyclotron frequency ratio , was calculated via Eq. 3. As shown in Fig. 3, 20 PI-ICR measurement data sets were collected for 97Tc+-97Mo+. In less than 5 minutes, a full scanning measurement (one cycle) of the magnetron phase, reduced cyclotron phase and center spot in sequence was completed for each ion species of the decay pair 97Tc+-97Mo+. In the analysis, we determined the position of each spot using the maximum likelihood method. After summing a few cycles to obtain reasonable counts for fitting, phase angles were calculated based on the determined positions of the phase spots as demonstrated in Fig. 2. These phase angles allowed us to deduce the cyclotron frequencies of each ion species. By linearly interpolating the of the daughter ion 97Mo+ (reference) to the time of the measurement of the parent ion 97Tc+ (ion of interest), the cyclotron frequency ratio was derived. To minimize potential cyclotron frequency shifts resulting from ion-ion interactions Kellerbauer et al. (2003); Roux et al. (2013), a median ion rate was constrained to 1-2 ions per bunch during the measurements. Bunches with fewer than five detected ions per bunch were considered during the data analysis. In the analysis, the count-rate related frequency shifts were not observed. The temporal fluctuation of the magnetic field /min Nesterenko et al. (2021) is taken in account in the final results. Here, represents the time interval between consecutive reference measurements. The contribution of magnetic field fluctuations to the final frequency ratio uncertainty was less than 10-10, as the parent-daughter measurements were interleaved with intervals less than 10 minutes. Additionally, the frequency shifts in the PI-ICR measurement caused by ion image distortions were ignored, as they were well below the statistical uncertainty. Furthermore, the fact that decay pair ions 97Tc+-97Mo+ are mass doublets, cancels out many systematic uncertainties in the cyclotron frequency ratio. The weighted mean ratio of all single frequency ratios was calculated, considering the inner and outer errors to deduce the Birge rato Birge (1932). The maximum of the inner and outer errors served as the weight for computing . Results of our analysis including all data sets, in comparison to literature values, are presented in Fig. 3 and Table 1. The final daughter-to-parent frequency ratio with its uncertainty is determined to be 1.000 003 598 4(23), resulting in a gs-to-gs value of 324.82(21) keV.
@@ IV Theoretical predictions
Two calculation methods are employed to predict the distribution of energy released in the decay, the transition half-life and their error spread, namely the atomic many-electron Dirac–Hartree–Fock–Slater (DHFS) self-consistent method and the Nuclear Shell Model (NSM) many-nucleon framework. For the atomic-structure calculations we made use of the RADIAL subroutine package Salvat and Fernández-Varea (2019), which also contains the DHFS.F code, while for the NSM we employed the NuShellX@MSU code Brown and Rae (2014).
@@
Channel | Interaction | Total half-life | |||||||
L1 allowed | jj45pnb | ||||||||
L1 allowed | glekpn | ||||||||
L1 forbidden | jj45pnb | ||||||||
L1 forbidden | glekpn |
Here we employ two corrections, namely the exchange and overlap correction represented by the factor and the shake-up and shake-off corrections introduced as the factor Mougeot (2018). The Coulomb amplitude is calculated for the initial atom and is represented by . The Cabibbo angle and the Fermi constant are present in . The relative occupancy of the shell is denoted as . The energy and momentum of the emitted neutrino are represented as and , with being the neutrino mass. A recent improvement Sevestrean et al. (2023) in the computation of the relaxation energy following the capture is the usage of total binding energy as , where indicates the energy of the final atom in its ground state, while represents the energy of the excited state of the final atom with a hole in the shell. The represents the intrinsic line-widths of Breit–Wigner resonances centered at and they were taken from CAMPBELL and PAPP (2001). The nuclear structure part is represented by the factor as:
@@
To compute the decay probability in the interval we integrated the distribution of the released energy:
The energy variable respects the relation . By using the narrow-width approximation, the total decay rate can be written as , with the partial decay constant defined as:
where .
In Table 3, we present the predicted half-lives and associated uncertainties for the decay of 97Tc to the excited state of 97Mo, corresponding to the experimental energy keV, across all relevant atomic shells. In the top two lines, we consider the L1 channel to be energetically possible (L1 allowed), while in the bottom two lines, we assume that the capture of L1 electrons is not energetically possible, but the capture from L2 is possible (L1 forbidden). As can be seen, there is a significant difference between the upper and lower error because the -value spans a considerable interval of values, and it appears in quadratic form in the half-life time formula.
@@
In Fig. 5 we present the normalized distributions of released energy within a shaded-in-blue 68 confidence band between the blue curves. Considering a Gaussian distribution for the Q-value 97.3 of cases permit the capture of an electron from the L1 shell (bottom plot). In 1.2 the L1 channel is energetically impossible but the capture from the L2 sub-shell is possible (top plot). The other 1.5 of cases were not presented here, due to their decreasing probability of occurrence. The black line represents the median values for each energy point in the graph. The interval was obtained as described in section IV.1. This analysis shows the high dependence of the normalized distribution on the value, especially in the interesting region at the end of the spectrum. The biggest impact in the error of comes from the uncertainty in the energy of the level of interest in the final nucleus. Thus a better assessment of the energy of this level is necessary in order to have a more well-defined distribution of energy at .
@@
The current measurement of the value with the associated error does not give a definite picture of the capture process. In order to investigate the implications of the associated uncertainties, we present in Fig. 6 a comparison between the normalized distribution of released energy as a function of the energy, , for the decay of 97Tc to excited state of 97Mo with keV, within 1 range of the experimental central value, and 163Ho to the ground state of 163Dy with keV. As can be seen in the enlarged scale in the inset of Fig. 6, the neutrino-mass-sensitive region provides a higher EC event rate for 97Tc for this Q value, making it comparable with the rate for 163Ho. Also, this inset shows exactly the impact of the neutrino mass on the shape of the spectrum and how it could be highlighted. The long half-life of about 4 My for 97Tc renders it a good candidate for future long-term neutrino-mass measurements.

### v2

@@ Abstract
A direct measurement of the ground-state-to-ground-state electron-capture decay () value of 97Tc has been conducted employing the high resolving power phase-imaging ion-cyclotron-resonance technique with the double Penning trap mass spectrometer JYFLTRAP. The resulting value for 97Tc is 324.82(21) keV, exhibiting a precision approximately 19 times higher than the value adopted in the newest Atomic Mass Evaluation (AME2020) and differing by 1.2. Furthermore, by combining this refined value with nuclear energy-level data for the decay-daughter 97Mo, a potential ultra-low Q-value transition 97Tc (9/2+, ground state) 97Mo∗ (320(1) keV), was detected. The ground-state-to-excited-state electron-capture decay value () of this transition was determined to be 4.8(10) keV, confirming it to be energetically allowed with a confidence level of exceeding 4. The captures of electrons occupying the L and higher shells for this transition are energetically allowed, giving a value of 2.0(10) keV for the closest distance of to the allowed binding energy of the L1 shell. To predict partial half-lives and energy-release distributions for this transition, the atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model have been employed. Dominant correction terms such as exchange and overlap corrections, as well as shake-up and shake-off effects, were included in the final results. Moreover, in the case of a possible allowed transition, the normalized distribution of released energy in the electron-capture decay of 97Tc to the excited 320-keV state of 97Mo, is compared with that of 163Ho, which is being used for electron-neutrino-mass determination. A pseudo-experiment technique was introduced to calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions.
@@ I Introduction
Neutrinos are electrically neutral fermions that interact only via the weak force and gravity, and they have an incredibly tiny rest mass compared to other elementary particles. Neutrinos do not participate in electromagnetic or strong interactions, thus allowing them to pass through matter unimpeded. Up to now, three neutrino flavors have been found: electron neutrino (), muon neutrino (), and tau neutrino (). For each neutrino, there also exists a corresponding antineutrino, which also has spin of 1/2 and no electric charge Petcov (2013). Although neutrinos were initially thought to be massless as assumed by the standard model (SM), the three neutrino mass eigenstates do not show a one-to-one correspondence to the three flavors, each flavor being a quantum superposition of all three mass states Lesgourgues and Pastor (2006). Neutrinos oscillate between flavors, making their detection even more intriguing Fukuda et al. (1998); SNO Collaboration (2002); Gerbino and Lattanzi (2018). Neutrino-oscillation experiments reveal that at least two neutrino-mass eigenstates have a non-zero rest mass, leading to a mass difference between the three mass eigenstates. Laboratory experiments and cosmological observations have provided experimental data on differences in the squares of the neutrino masses, an upper limit on their sum, and an upper limit on the mass of the electron (anti)neutrino. Among these experiments, the most direct and model-independent way to pin down the exact mass value of the electron antineutrino (m) is to measure the electron energy spectrum of single decay, e.g., the Karlsruhe Tritium Neutrino (KATRIN) experiment Drexlin et al. (2013); Aker et al. (2019, 2022), which has recently pushed the absolute electron antineutrino mass scale boundary to 0.45 eV/c2 at the 90% confidence level (CL) KATRIN Collaboration et al. (2025). KATRIN will finally set the m upper limit with a sensitivity of 0.2 eV/c2 at 90% CL. Other single--decay experiments, Project 8 and PTOLEMY project, using cyclotron radiation emission spectroscopy and the development of atomic tritium sources, will eventually allow a sensitivity of tritium-based neutrino-mass experiments to m beyond the design sensitivity of KATRIN Esfahani et al. (2017); Betti et al. (2019) down to 0.04 eV/c2. An alternative method, employing electron capture in 163Ho (ECHo), an experiment to calorimetrically measure the electron-capture (EC) deexcitation spectrum Gastaldo et al. (2014, 2017); Velte et al. (2019), has set a current limit of 150 eV/c2 for the electron neutrino mass Velte et al. (2019).
@@
In single--decay experiments aimed at determining the mass of the electron antineutrino, a value of the decay as small as possible is sought. A smaller value leads to a higher fraction of effective decays within a given energy interval near the endpoint region. Currently, the running direct (anti)neutrino-mass-determination experiments rely exclusively on ground-state-to-ground-state (gs-to-gs) decay cases, such as 3H ( decay) and 163Ho (EC). However, ongoing intensive searches are actively exploring isotopes having low -value ground-state-to-excited-state (gs-to-es) /EC decays Cattadori et al. (2007); Haaranen and Suhonen (2013); Suhonen (2014); Ejiri et al. (2019) using high-precision Penning-trap mass spectrometry (PTMS) including JYFLTRAP, LEBIT, CPT, ISOTRAP, and SHIPTRAP Penning traps Sandler et al. (2019); Karthein et al. (2019); De Roubin et al. (2020); Ge et al. (2021a, b, 2022a); Eronen et al. (2022); Ge et al. (2022a, b); Ramalho et al. (2022); Gamage et al. (2022); Keblbeck et al. (2023); Ge et al. (2023); Valverde et al. (2024); Ge et al. (2024a, b). PTMS is the most precise and accurate technique for determining atomic masses and values. It remains the only direct method capable of measuring decay values to sub-keV precision or better, allowing verification of potential ultra-low (¡ 1 keV) -value transitions.
@@ II Experimental method
To minimize any shifts in the ratio of the 97Tc+-97Mo+ pair, the positions of the phase spots for the magnetron and reduced cyclotron motions were carefully chosen to maintain the angle within a few degrees. This was done to account for the conversion of the reduced cyclotron motion to the magnetron motion and the potential distortion of the ion-motion projection onto the detector, ensuring that the shift remained well below 10-10 Eliseev et al. (2014). Figure 2 shows a representative measurement with ”reduced cyclotron” and ”magnetron” phase spots relative to the center spot. During the measurements, the excitation delay of was systematically scanned over one magnetron period, while the extraction delay varied over one cyclotron period, accounting for any residual motion that might have affected the spots. The interleaved measurements of for 97Tc+-97Mo+ ions had a total data accumulation time of approximately 4.9 hours.
@@
| Final state |
| 97Mo () | 320.0(40) | 324.82(21) | 0.0(41) | 4.8(10) | 320.0(10) | 4.8 | 2.0(10) | 2.2(10) | 4.3(10) | ||||||||||||||||
@@ III Results and discussion
From the measurements of the cyclotron frequency ratio , was calculated via Eq. 3. As shown in Fig. 3, 20 PI-ICR measurement data sets were collected for 97Tc+-97Mo+. In less than 5 minutes, a full scanning measurement (one cycle) of the magnetron phase, reduced cyclotron phase and center spot in sequence was completed for each ion species of the decay pair 97Tc+-97Mo+. In the analysis, we determined the position of each spot using the maximum likelihood method. After summing a few cycles to obtain reasonable counts for fitting, phase angles were calculated based on the determined positions of the phase spots as demonstrated in Fig. 2. These phase angles allowed us to deduce the cyclotron frequencies of each ion species. By linearly interpolating the of the daughter ion 97Mo+ (reference) to the time of the measurement of the parent ion 97Tc+ (ion of interest), the cyclotron frequency ratio was derived. To minimize potential cyclotron frequency shifts resulting from ion-ion interactions Kellerbauer et al. (2003); Roux et al. (2013), a median ion rate was constrained to 1-2 ions per bunch during the measurements. Bunches with fewer than five detected ions per bunch were considered during the data analysis. In the analysis, the count-rate related frequency shifts were not observed. The temporal fluctuation of the magnetic field /min Nesterenko et al. (2021) is taken in account in the final results. Here, represents the time interval between consecutive reference measurements. The contribution of magnetic field fluctuations to the final frequency ratio uncertainty was less than 10-10, as the parent-daughter measurements were interleaved with intervals less than 10 minutes. Additionally, the frequency shifts in the PI-ICR measurement caused by ion image distortions were ignored, as they were well below the statistical uncertainty. Furthermore, the fact that decay pair ions 97Tc+-97Mo+ are mass doublets, cancels out many systematic uncertainties in the cyclotron frequency ratio. The weighted mean ratio of all single frequency ratios was calculated, considering the inner and outer errors to deduce the Birge ratio Birge (1932). The maximum of the inner and outer errors served as the weight for computing . Results of our analysis including all data sets, in comparison to literature values, are presented in Fig. 3 and Table 1. The final daughter-to-parent frequency ratio with its uncertainty is determined to be 1.000 003 598 4(23), resulting in a gs-to-gs value of 324.82(21) keV.
@@ IV Theoretical predictions
Two calculation methods are employed to predict the distribution of energy released in the decay, the transition half-life and their error spread, namely the atomic many-electron Dirac–Hartree–Fock–Slater (DHFS) self-consistent method and the Nuclear Shell Model (NSM) many-nucleon framework. For the atomic-structure calculations we made use of the RADIAL subroutine package Salvat and Fernández-Varea (2019), which also contains the DHFS.F code, while for the NSM we employed the NuShellXMSU code Brown and Rae (2014).
@@
| LCS | Decay | Total half-life | ||||||||
| type | ||||||||||
| L1 | allowed | |||||||||
| L2 | allowed | |||||||||
| Total half-life | ||||||||||
| L1 | 2nd UF | |||||||||
| L2 | 2nd UF | |||||||||
| Total half-life | ||||||||||
| L1 | 4th NUF | |||||||||
| L2 | 4th NUF |
Here we employ two corrections, namely the exchange and overlap correction represented by the factor and the shake-up and shake-off corrections introduced as the factor Mougeot (2018). The Coulomb amplitude is calculated for the initial atom and is represented by . The Cabibbo angle and the Fermi constant are present in . The relative occupancy of the shell is denoted as . The energy and momentum of the emitted neutrino are represented as and , with being the neutrino mass. A recent improvement Sevestrean et al. (2023) in the computation of the relaxation energy following the capture is the usage of total binding energy as , where indicates the energy of the final atom in its ground state, while represents the energy of the excited state of the final atom with a hole in the shell. The represents the intrinsic line-widths of Breit–Wigner resonances centered at and they were taken from CAMPBELL and PAPP (2001).
The spin of the 320-keV excited state in 97Mo is uncertain, ranging within a range of . This means that the decay transition from the ground state of 97Tc to the 320-keV state could be allowed ( and ), second-forbidden non-unique (2nd NUF) in the case of , second-forbidden unique (2nd UF) in the case of and fourth-forbidden non-unique (4th NUF) in the case of . In the case of an allowed transition, the nuclear structure part is represented by the factor
@@
To compute the decay probability we employ the formalism described in Bambynek et al. (1977):
where . For allowed transition we used the simplified approach for the factor in Eq. 5, while for the unique and non-unique forbidden transitions, is defined as:
The and factors are defined in Bambynek et al. (1977).
In Table 3, we present the predicted half-lives and associated uncertainties for the decay of 97Tc to the excited state of 97Mo, corresponding to the experimental energy keV, across all relevant atomic shells. Due to the uncertainty of the angular momentum of the final state, we considered multiple possible values. By computing the energy levels using the jj45pnb interaction, we obtained a few possible values for J in the vicinity of 320 keV, namely 7/2, 3/2, and 1/2. Thus, we have the possible decay types allowed, 2nd unique forbidden, and 4th non-unique forbidden. The matrix elements needed for the half-life were computed with the same jj45pnb interaction. We also used the glekpn interaction for the allowed transition; the computed half-life was an order of magnitude smaller. Comparing the half-life for the excited state transition with the total half-life measured, years NND (2021), one observes that for the allowed transition, the computed half-life is in tension with the measured one. As such, this is most likely due to the assumed angular momentum of the excited final state, which thus should be verified by new spectroscopic measurements. Considering the uncertainty in the value, the capture from the L1 shell is uncertain but we retain this possibility in our computations. However, the capture from the L2 shell is certainly possible. As can be seen, there is a significant difference between the upper and lower error because the -value spans a considerable interval of values, and it appears at various powers in the half-life formula.
@@
In Fig. 5 we present the normalized distributions of released energy within a shaded-in-blue 68 confidence band between the blue curves assuming an allowed transition. Considering a Gaussian distribution for the Q-value, 97.3 of cases permit the capture of an electron from the L1 shell (bottom plot). In 1.2 the L1 channel is energetically impossible but the capture from the L2 sub-shell is possible (top plot). The other 1.5 of cases were not presented here, due to their decreasing probability of occurrence. The black line represents the median values for each energy point in the graph. The interval was obtained as described in section IV.1. This analysis shows the high dependence of the normalized distribution on the value of , especially in the interesting region at the end of the spectrum. The greatest impact on the error of comes from the uncertainty in the energy of the level of interest in the final nucleus. Thus a better assessment of the energy of this level is necessary in order to have a more well-defined distribution of energy at .
@@
The current measurement of the value of with the associated error does not give a definite picture of the capture process. In order to investigate the implications of the associated uncertainties in the case of an allowed transition, we present in Fig. 6 a comparison between the normalized distribution of the released energy as a function of the energy, , for the decay of 97Tc to the excited state of 97Mo with keV, within 1 range of the experimental central value, and 163Ho to the ground state of 163Dy with keV. As can be seen in the enlarged scale in the inset of Fig. 6, the neutrino-mass-sensitive region provides a higher EC event rate for 97Tc for this Q value, making it comparable with the rate for 163Ho. Also, this inset shows exactly the impact of the neutrino mass on the shape of the spectrum and how it could be highlighted. A strong 97Tc source can be produced via the reaction with 19 MeV protons on an enriched 97Mo target (cross-section 206 mbarn from theoretical calculation with TALYS Koning, Arjan et al. (2023)). With a 10 mA beam and 1 month irradiation, we estimate producing atoms, sufficient for a neutrino mass experiment requiring sub-keV sensitivity over one year if the allowed transition can be confirmed. The long half-life of about 4 My for 97Tc renders it a good candidate for future long-term neutrino-mass measurements if the considered transition is of the allowed type. In case the transition is of forbidden type, it is not relevant for neutrino-mass measurements as witnessed by the very long half-lives of the transition in Table 3.